OpenStudy (anonymous):

f(x) = (15x^2 -7x -2)/(x^2 -4) find the x intercept?

OpenStudy (anonymous):

Take y = 0 and cross multiply, we are left with 15x^2 - 7x - 2 which is a quadratic equation. 15x^2 - 7x - 2 Solve by factoring or completing the square or determinant method.

OpenStudy (anonymous):

i tried to factor and got it completely wrong, i got x-10 and x+3 @Fall12-13

OpenStudy (anonymous):

Nah, try determinant method.

OpenStudy (anonymous):

hows that

OpenStudy (anonymous):

id rather do it by factoring, thats how my teacher wants it to be done

OpenStudy (anonymous):

$$f(x)=\frac{15x^2-7x-2}{x^2-4}\\f(x)=0\\\frac{15x^2-7x-2}{x^2-4}=0\\15x^2-7x-2=0\\15x^2+3x-10x-2=0\\3x(5x+1)-2(5x+1)=0\\(3x-2)(5x+1)=0\\3x-2=0\ \lor\ 5x+1=0\\x\in\left\{\frac23,-\frac15\right\}$$

OpenStudy (anonymous):

@oldrin.bataku thanks bro, i got that i just wrote the answer in 2 different sets